OpenStudy (anonymous):

Help plz: In a right triangle, the length c of the hypotenuse is 14 and the length b of one of the legs is 8. Find the length a of the other leg to the nearest hundredth. A:21.23 B:19.88 C:16.12 D:11.49

OpenStudy (anonymous):

By the Pythagorean theorem, if the hypotenuse is C and the legs are A and B, then C^2 = A^2 + B^2 Plug in the hypotenuse and the leg you're given and solve for the other leg.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What are you confused about?

OpenStudy (anonymous):

I'll give you an example. If I know that the hypotenuse of a triangle is 5 and one leg is 3, then I can plug into the Pythagorean theorem like this: C^2 = A^2 + B^2 5^2 = 3^2 + B^2 now solve for B. 25 = 9 + B^2 25-9 = B^2 16 = B^2 sqrt(16) = B 4 = B.

OpenStudy (anonymous):

Sorry. I teach math. I don't give answers.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Happy to explain if you have any questions on how to do it.

OpenStudy (anonymous):

thinking

OpenStudy (anonymous):

i would say c because it goes best with what you said above

OpenStudy (anonymous):

Thanks for the help